# On the applicability of Kolmogorov’s theory of probability

### v1

@@ Abstract
It is a common view that von Neumann laid the foundations of a “non-commutative probability theory” with his axiomatization of quantum mechanics (QM). As such, it is regarded a generalization of the “classical probability theory” due to Kolmogorov. Outside of quantum physics, however, Kolmogorov’s axioms enjoy universal applicability. This raises the question of whether quantum physics indeed requires such a generalization of our conception of probability or if von Neumann’s axiomatization of QM was contingent on the absence of a general theory of probability in the 1920s.
@@ 3 Physical considerations
Second, the most direct result on the applicability of Kolmogorov’s axioms to quantum phenomena, which I am aware of, is the BCHSH inequality [clauserProposedExperimentTest1969, morettiFundamentalMathematicalStructures2019]. This constitutes a generalization of Bell’s theorem [bellEinsteinPodolskyRosen1964], that overcomes the “unrealistic restriction that for some pair of parameters […] there is perfect correlation” [clauserProposedExperimentTest1969]. If one follows the statement that those inequalities are empirically violated, then the random variable in question may not factorize. Whether the approach described here supports this kind of “non-locality” needs to be investigated in the context of a respective dynamical theory [takabayasiDescriptionPauliMatter1957, janossyHydrodynamicalModelWave1966]. $\diamondsuit$
@@ 6 A hybrid theory
We consider the simplest quantum-mechanical model of the hydrogen atom, i.e. $N=1$ and $V(\vec{r})=\text{e}/(4\pi\varepsilon_{0}\left\lvert\vec{r}\right\rvert)$. Using physics conventions, we denote the electron mass by $\mu$ and each stationary state basis function by $\Psi_{nlm}$; $n$, $l$, and $m$ are appropriately chosen integers.
@@ 7 Conclusion
Acknowledgments I would like to thank Bill Poirier, Mike Scherfner, and Max von Renesse for their indirect support of my research efforts. I also thank Jamal Berakdar for helpful comments and Mario Flory for inviting me to a conference, where I held a similarly titled talk. Furthermore, I am indebted to Anhalt UAS for enabling my participation therein.

### v2

@@ Abstract
It is a common view that von Neumann laid the foundations of a “quantum probability theory” with his axiomatization of quantum mechanics (QM). As such, it is regarded a generalization of the “classical probability theory” due to Kolmogorov. Outside of quantum physics, however, Kolmogorov’s axioms enjoy universal applicability. This raises the question of whether quantum physics indeed requires such a generalization of our conception of probability or if von Neumann’s axiomatization of QM was contingent on the absence of a general theory of probability in the 1920s.
@@ 3 Physical considerations
Second, the most direct result on the applicability of Kolmogorov’s axioms to quantum phenomena, which I am aware of, is the BCHSH inequality [clauserProposedExperimentTest1969, morettiFundamentalMathematicalStructures2019]. This constitutes a generalization of Bell’s theorem [bellEinsteinPodolskyRosen1964], that overcomes the “unrealistic restriction that for some pair of parameters […] there is perfect correlation” [clauserProposedExperimentTest1969]. If one follows the statement that those inequalities are empirically violated, then, by symmetry considerations, for any unit vectors $\vec{a}$, $\vec{b}\in\mathbb{R}^{3}$ the random variable $X(\vec{a},\vec{b},\,.\,)$ in question must be
$X(\vec{a},\vec{b},\,.\,)=A(\vec{a},\vec{b},\,.\,)\,B(\vec{a},\vec{b},\,.\,)$ (3.2)
instead of $A(\vec{a},\,.\,)\,B(\vec{b},\,.\,)$. Whether the approach described here supports this kind of “(Bell) non-locality” needs to be investigated in the context of a respective dynamical theory [takabayasiDescriptionPauliMatter1957, janossyHydrodynamicalModelWave1966]. $\diamondsuit$
@@ 6 A hybrid theory
We consider the simplest quantum-mechanical model of the hydrogen atom, i.e. $N=1$ and $V(\vec{r})=-\text{e}^{2}/(4\pi\varepsilon_{0}\left\lvert\vec{r}\right\rvert)$. Using physics conventions, we denote the electron mass by $\mu$ and each stationary state basis function by $\Psi_{nlm}$; $n$, $l$, and $m$ are appropriately chosen integers.
@@ 7 Conclusion
Acknowledgments I would like to thank Bill Poirier, Mike Scherfner, and Max von Renesse for their indirect support of my research efforts. I also thank Jamal Berakdar for helpful comments, Wolfgang Paul for helpful discussion, and Mario Flory for inviting me to a conference, where I held a similarly titled talk. Furthermore, I am indebted to Anhalt UAS for enabling my participation therein.